# On-the-fly training of polynomial machine learning potentials in computing lattice thermal conductivity

### v1

@@ Abstract
The application of first-principles calculations for predicting lattice thermal conductivity (LTC) in crystalline materials, in conjunction with the linearized phonon Boltzmann equation, has gained increasing popularity. In this calculation, the determination of force constants through first-principles calculations is critical for accurate LTC predictions. For material exploration, performing first-principles LTC calculations in a high-throughput manner is now expected, although it requires significant computational resources. To reduce computational demands while preserving accuracy moderately, we integrated polynomial machine learning potentials on-the-fly during the first-principles LTC calculations. This paper presents a systematic approach to first-principles LTC calculations. We designed and optimized an efficient workflow that integrates multiple modular software packages. We applied this approach to calculate LTCs for 103 compounds of the wurtzite, zincblende, and rocksalt types to evaluate the performance of the polynomial machine learning potentials in LTC calculations. We demonstrate a significant reduction in the computational resources required for the LTC predictions, while maintaining reasonable accuracy.
@@ I Introduction
Calculations of lattice thermal conductivity (LTC) based on first-principles calculations and the linearized phonon Boltzmann equation Peierls (1929, 2001); Ziman (1960); Srivastava (1990) have become increasingly popular in recent years. This is because sufficiently accurate LTC values can be systematically predicted for a wide variety of crystals using available computer simulation packages.Tadano et al. (2014); Li et al. (2014); Carrete et al. (2017); Chernatynskiy and Phillpot (2015); Togo et al. (2015) These computational tools are expected to be applied in materials discovery within a high-throughput calculation environment. However, since first-principles LTC calculations are still computationally intensive, there is a need for the development of methodologies to reduce the computational demands.
@@
The accuracy of predicting LTCs relies on the use of first-principles calculations to obtain the displacement-force dataset. However, this approach is computationally intensive. In order to achieve precise LTC predictions with lower computational resources, compressive sensing force constants calculation methods were developed, as reported in Refs. Zhou et al., 2014; Tadano and Tsuneyuki, 2015. These methods employ regularized linear regression techniques to eliminate certain tensor elements of the supercell force constants, thereby reducing the required size of the displacement-force dataset.
In this study, we introduce another approach to reduce the computational demands of first-principles LTC calculations. We incorporate polynomial machine learning potentials (MLPs) Seko et al. (2019); Seko (2023) into an intermediate stage of the LTC calculation process. The polynomial MLPs are trained using a small dataset of displacement-force pairs and energies derived from first-principles calculations. Subsequently, the polynomial MLPs generate a large displacement-force dataset to calculate supercell force constants with significantly lower computational demands than those required by first-principles calculations.
@@
In this study, a systematic calculation of LTCs at 300 K was performed for the same set of the 103 compounds of wurtzite, zincblende, and rocksalt types reported in Ref. Togo, 2023. The computational workflow-design and details that we reached after struggling to achieve accuracy are presented in Secs. II and III. While most of the theoretical and methodological background is covered in the referenced articles, those relevant to this paper are described in Sec. III. The results of the LTC calculations are summarized in Sec. IV.
@@ II Design of computational workflow
As shown in Fig. 1, our study utilized a specific approach for LTC calculations. This approach consists of specialized modules for each calculation step. These modules are interconnected through data communication, including computer-network file transfers and application program interfaces (APIs).
@@
In the LTC calculation process, we generate two distinct sets of supercells with random displacements of atoms. This occurs in steps (a) and (d), as illustrated in Fig. 1. In step (b), we conduct energy and force calculations for the first set of supercells using first-principles calculations. This is the most computationally demanding step. The outputs of step (b) form a dataset used to train polynomial MLPs in step (c). In step (e), we calculate forces for the second set of supercells using the trained polynomial MLPs. In step (f), third-order supercell force constants are computed using the displacement-force dataset from step (e). Finally in step (g), LTC is calculated from the supercell force constants. Use of crystal symmetry is important in steps (f) and (g) for the computational efficiency and numerical accuracy. The computational demand from step (c) to the end is negligible compared to step (b).
@@
Our computational workflow is specifically designed to optimize high-throughput LTC calculations, balancing efficiency and convenience. This workflow is divided into two main parts: dataset preparation and calculations using this dataset. The dataset preparation stage involves a bunch of energy and force calculations using first-principles calculations, which are normally distributed over computer nodes to conduct the calculations in parallel. After completing these calculations, we extract the necessary data from the output files of the first-principles calculations. This data is then saved for transfer via computer-network communication. In the subsequent steps, calculations are performed on a single computer. Depending on the operational requirements, data transfer between the modules is facilitated either through APIs or via the local computer file system for ease of use.
@@
Figure 1: Schematic illustration of workflow employed in this study for LTC calculations. Each calculation step is represented by a box, within which the name of the software package and the type of calculation are described. The arrows roughly indicate flow of data. The data were passed via APIs or computer files. The steps are as follows: Step (a): Generation of a modest number of supercells with random displacements. Step (b): Calculation of energies and forces in the supercells generated in step (a) by first-principles calculations. The submission of a large number of computational jobs is automated using a workflow system. Step (c): Training of polynomial MLPs using the energies and forces in the supercells obtained in step (b). Step (d): Generation of a large number of supercells with random displacements. Step (e): Calculation of forces in the supercells generated in step (d) using the trained polynomial MLPs. Step (f): Calculation of force constants from the displacement-force dataset calculated in step (e). Step (g): Calculation of LTC using the supercell force constants obtained in step (f). In the conventional approach, the dataset obtained in step (b) is directly used in step (f).
@@ III.1 LTC calculation
LTCs were computed by solving the Peierls-Boltzmann equation within the relaxation time approximation (RTA)Peierls (1929, 2001); Allen and Perebeinos (2018) using the phono3py code.Togo et al. (2015, 2023) To simplify the methodological investigation in this study, we considered only phonon-phonon scattering for determining the phonon relaxation times. Imaginary parts of phonon self-energies corresponding to the bubble diagram were calculated from supercell force constants, and their reciprocals were used as the relaxation times. Computational details on the calculation of the supercell force constants are provided in the subsequent sections. Phonon group velocities, mode heat capacities, frequencies, and eigenvectors were obtained from dynamical matrices constructed using second-order supercell force constants. Additionally, a non-analytical term correction Pick et al. (1970); Gonze et al. (1994); Gonze and Lee (1997) was applied to the dynamical matrices to account for long-range dipole-dipole interactions in the harmonic phonon calculation. For the wurtzite-type compounds, reciprocal spaces were sampled using a $19\times 19\times 10$ mesh, while a $19\times 19\times 19$ mesh was employed for the zincblende-type and rocksalt-type compounds.
@@ III.2 Supercell force constants calculation
The linear regression method was employed to calculate supercell force constants from the displacement-force dataset. In this method, forces acting on atoms, which were randomly displaced from their equilibrium positions in supercells, were computed either using the polynomial MLPs in our current approach or through first-principles calculations in the conventional approach. Tensor elements of supercell force constants were projected onto the subspace defined by symmetry projection operators of totally symmetric irreducible representations of the space group, index permutation, and translational invariance. This process is implemented in the symfc code.Seko and Togo The crystallographic symmetries were determined using the spglib code.Togo and Tanaka (2018) Various techniques were developed to make this computational process efficient, but these technical details are beyond the scope of this paper and will be described elsewhere.
For the zincblende- and rocksalt-type compounds, we utilized supercells with $2\times 2\times 2$ and $4\times 4\times 4$ expansions of the conventional unit cells to calculate third-order and second-order supercell force constants, respectively. In the case of the wurtzite-type compounds, supercells with $3\times 3\times 2$ and $5\times 5\times 3$ expansions of the unit cells were employed for third-order and second-order force constants calculations, respectively.
To compute the second-order supercell force constants, we employed the finite difference method as implemented in the phonopy code.Togo et al. (2023); Togo (2023) We used the same displacement-force datasets as those in Ref. Togo, 2023, where the forces in these datasets had been computed through first-principles calculations. The number of supercells in the datasets were six, two, and two for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively.
Third-order supercell force constants were calculated from the displacement-force datasets using the symfc code.Seko and Togo The numbers of symmetrically independent force constant elements were 7752, 1536, and 758 for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively. These values were determined based on the forces acting on atoms, which were inferred using the polynomial MLPs implemented in the pypolymlp code.Seko (2023)
@@ III.3 Polynomial MLPs
The polynomial MLPs were trained using the dataset composed of forces and displacements of atoms and energies in supercells. These energies and forces were computed through first-principles calculations. The performance of the polynomial MLPs in the LTC calculation via the third-order force constants is discussed in Section IV.
For the 103 compounds, we trained the polynomial MLPs using the pypolymlp code.Seko (2023) In this training, Gaussian-type radial functions were employed, and the functional form $f_{n}(r)$ is given as
@@
where $r$ represents the distance from the center of each atom, and $r_{\text{c}}$ is the cutoff distance. $\beta_{n}$ and $r_{n}$ are the parameters, respectively. Radial functions with $r_{\text{c}}=8.0$ and $\beta_{n}=1.0$${}^{-2}$ and $r_{n}=(n-1)(r_{\text{c}}-1.0)/11$ for $n=1,\ldots,12$ were used. We considered polynomial invariants up to third order characterizing neighboring atomic density based on spherical harmonics with the maximum angular numbers of spherical harmonics $l_{\text{max}}^{(2)}=l_{\text{max}}^{(3)}=8$. The polynomial models were then constructed by the polynomial functions of the pair invariants and linear polynomial function of the polynomial invariants. We considered polynomial functions up to second order. The model coefficients were estimated from electronic total energies and forces by the linear ridge regression method.
@@ III.4 First-principles calculation
For the first-principles calculations, we employed the plane-wave basis projector augmented wave (PAW) method Blöchl (1994) within the framework of DFT as implemented in the VASP code.Kresse (1995); Kresse and Furthmüller (1996); Kresse and Joubert (1999) The generalized gradient approximation (GGA) of Perdew, Burke, and Ernzerhof revised for solids (PBEsol) Perdew et al. (2008) was used as the exchange correlation potential. To ensure high numerical accuracy in computing atomic forces, the projection operators were applied in reciprocal spaces and additional support grids were employed for the evaluation of the augmentation charges. Static dielectric constants and Born effective charges were calculated with the conventional unit cells from density functional perturbation theory (DFPT) as implemented in the VASP code.Gajdoš et al. (2006); Wu et al. (2005)
@@ III.5 Automation of dataset preparation
For the systematic calculations of energies and forces in supercells using first-principles calculations, we utilized the AiiDA environment Pizzi et al. (2016) in conjunction with the AiiDA-VASP plugin.Aii The advantage of using the workflow automation system was not only the automation of submitting calculation jobs to high-performance computers, but also the automated data storing of the calculation results in a database systematically. The computed data, stored within the AiiDA database, could be conveniently accessed through the Python programming language. By writing a concise Python script, we were able to extract supercell energies, forces, and displacements from the AiiDA database on demand and convert this data into the structured format required for immediate use by the phono3py code.Togo et al. (2015, 2023)
@@ III.6 Parameters for 103 binary compounds
33 compounds for the wurtzite- and zincblende-type and 37 compounds for the rocksalt-type were used to evaluate the LCT calculation approach proposed in this study, and their chemical compositions are listed in Tables 1 and 2. Crystal structures of the wurtzite and zincblende types are similar, though their stacking orders are different, much like the relationship between face-centered-cubic and hexagonal-close-packed structure types. Since it is of interest to explore their similarities and differences in calculations, as also studied in Ref. Togo et al., 2015, the compounds with the same chemical compositions for the wurtzite and zincblende types were calculated. The tables also provide information on lattice parameters, the choices of PAW datasets from the VASP package, and electronic total energies of the elements that were subtracted from the total energies of the compounds used to train the polynomial MLPs.
Table 1: Lattice parameters, names of the VASP PAW-PBE datasets, electronic total energies of the atoms used in this study for 33 wurtzite- and zincblende-type compounds. w-$a$, w-$c$, and z-$a$ denote the lattice perameters a and c of the wurtzite-type compounds, and a of the zincblende-type compounds, respectively.
@@ IV.1 Choice of displacements and number of supercells
For strongly anharmonic crystals, self-consistent phonon methods, which were not considered in this study, may have to be employed to obtain physically more meaningful results. In such cases, the use of the polynomial MLPs can also be beneficial to accelerate the LTC calculations.
@@ IV.2 Calculated LTCs
LTC values at 300 K predicted by the conventional approach, which directly uses the displacement-force dataset obtained through first-principles calculations to fit third-order supercell force constants, are depicted by the horizontal dotted lines. The third-order supercell force constants were computed by the linear regression method as implemented in the symfc code Seko and Togo from the first datasets with 100 supercells and 0.03 Å random directional displacements, which were those prepared for training the polynomial MLPs, as explained in Sec. IV.1. LTC values predicted using the polynomial MLPs for the wurtzite-type compounds tend to be larger than those represented by the horizontal lines. This may indicate that achieving better convergence in LTC values in the conventional approach might require more than 100 supercells. Although we have not investigated systematically the convergence of the LTC values in the conventional approach, we consider that the LTC values of the zincblende-type compounds can converge with smaller datasets than those of the wurtzite-type compounds due to the high symmetry of zincblende-type crystals compared to wurtzite-type crystals.
Figure 2: Filled circles show LTCs ($\kappa$) of the 33 wurtzite-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The LTC values are the averages of the diagonal elements, i.e., $(2\kappa_{xx}+\kappa_{zz})/3$. The horizontal dotted lines depict the LTC values calculated in the conventional approach from the datasets of 100 supercells without using the polynomial MLPs.
Figure 3: Filled circles show LTCs of the 33 zincblende-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The horizontal dotted lines depict the LTC values calculated in the conventional approach from the datasets of 100 supercells without using the polynomial MLPs.
Figure 4: Filled circles show LTCs of the 37 rocksalt-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The horizontal dotted lines depict the LTC values calculated in the conventional approach from the datasets of 100 supercells without using the polynomial MLPs.

### v3

@@ Abstract
The application of first-principles calculations for predicting lattice thermal conductivity (LTC) in crystalline materials, in conjunction with the linearized phonon Boltzmann equation, has gained increasing popularity. In this calculation, the determination of force constants through first-principles calculations is critical for accurate LTC predictions. For material exploration, performing first-principles LTC calculations in a high-throughput manner is now expected, although it requires significant computational resources. To reduce computational demands, we integrated polynomial machine learning potentials on-the-fly during the first-principles LTC calculations. This paper presents a systematic approach to first-principles LTC calculations. We designed and optimized an efficient workflow that integrates multiple modular software packages. We applied this approach to calculate LTCs for 103 compounds of the wurtzite, zincblende, and rocksalt types to evaluate the performance of the polynomial machine learning potentials in LTC calculations. We demonstrate a significant reduction in the computational resources required for the LTC predictions.
@@ I Introduction
Calculations of lattice thermal conductivity (LTC) based on first-principles calculations and the linearized phonon Boltzmann equation [1, 2, 3, 4] have become increasingly popular in recent years. This is because sufficiently accurate LTC values can be systematically predicted for a wide variety of crystals using available computer simulation packages.[5, 6, 7, 8, 9, 10] These computational tools are expected to be applied in materials discovery within a high-throughput calculation environment. However, since first-principles LTC calculations are still computationally intensive, there is a need for the development of methodologies to reduce the computational demands.
@@
The accuracy of predicting LTCs relies on the use of first-principles calculations to obtain the displacement-force dataset. However, this approach is computationally intensive. In order to achieve precise LTC predictions with lower computational resources, compressive sensing force constants calculation methods were developed, as reported in Refs. 11, 12. These methods employ regularized linear regression techniques to eliminate certain tensor elements of the supercell force constants, thereby reducing the required size of the displacement-force dataset.
In this study, we introduce another approach to reduce the computational demands of first-principles LTC calculations. We incorporate polynomial machine learning potentials (MLPs) [13, 14] into an intermediate stage of the LTC calculation process. The polynomial MLPs are trained using a small dataset of displacement-force pairs and energies derived from first-principles calculations. Subsequently, the polynomial MLPs generate a large displacement-force dataset to calculate supercell force constants with significantly lower computational demands than those required by first-principles calculations. Compared to the compressive sensing approach, this approach may more efficiently represent energy surface of crystal potential. However, the workflow for performing the first-principles LTC calculation with this approach is more complex. One objective of this study is to encapsulate this intricate workflow into a software package, complete with a set of well-optimized default parameters. Developing this software package requires a thorough understanding of the MLP code details. Moreover, efficient estimations of polynomial MLPs can be accomplished using linear regressions supported by powerful libraries for linear algebra. Therefore, we have utilized the polynomial MLP code developed by one of the authors.
@@
In this study, a systematic calculation of LTCs at 300 K was performed for the same set of the 103 compounds of wurtzite, zincblende, and rocksalt types reported in Ref. 15. The computational workflow-design and details are presented in Secs. II and III. While most of the theoretical and methodological background is covered in the referenced articles, those relevant to this paper are described in Sec. III. The results of the LTC calculations are summarized in Sec. IV.
@@ II Design of computational workflow
As shown in Fig. 1, our study utilized a specific approach for LTC calculations. This approach consists of specialized modules for each calculation step. These modules are interconnected through a local file system and data communication, including computer-network file transfers and application program interfaces (APIs).
@@
In the LTC calculation process, we generate two distinct sets of supercells with random displacements of atoms. This occurs in steps (a) and (d), as illustrated in Fig. 1. In this study, we displaced atoms by a constant distance in random directions. In step (b), we conduct energy and force calculations for the first set of supercells using first-principles calculations. This is the most computationally demanding step. The outputs of step (b) form a dataset used to train polynomial MLPs in step (c). In step (e), we calculate forces for the second set of supercells using the trained polynomial MLPs. In step (f), third-order supercell force constants are computed using the displacement-force dataset from step (e). Finally in step (g), LTC is calculated from the supercell force constants. Use of crystal symmetry is important in steps (f) and (g) for the computational efficiency and numerical accuracy. The computational demand from step (c) to the end is negligible compared to step (b).
@@
Our computational workflow is specifically designed to optimize high-throughput LTC calculations, balancing efficiency and convenience. This workflow is divided into two main parts: dataset preparation and calculations using this dataset. The dataset preparation stage involves a set of energy and force calculations using first-principles calculations, which are normally distributed over computer nodes to conduct the calculations in parallel. After completing these calculations, we extract the necessary data from the output files of the first-principles calculations. This data is then saved for transfer via computer-network communication. In the subsequent steps, calculations are performed on a single computer. Depending on the operational requirements, data transfer between the modules is facilitated either through APIs or via the local computer file system for ease of use.
@@
Figure 1: Schematic illustration of workflow employed in this study for LTC calculations. Each calculation step is represented by a box, within which the name of the software package and the type of calculation are described. The arrows roughly indicate flow of data. The data were passed via APIs or computer files. The steps are as follows: Step (a): Generation of a modest number of supercells with random displacements. Step (b): Calculation of energies and forces in the supercells generated in step (a) by first-principles calculations. The submission of a large number of computational jobs is automated using a workflow system. Step (c): Training of polynomial MLPs using the energies and forces in the supercells obtained in step (b). Step (d): Generation of a large number of supercells with random displacements. Step (e): Calculation of forces in the supercells generated in step (d) using the trained polynomial MLPs. Step (f): Calculation of force constants from the displacement-force dataset calculated in step (e). Step (g): Calculation of LTC using the supercell force constants obtained in step (f). This study introduces steps (c), (d), and (e) to the workflow. Conventionally, the dataset from step (b) is used directly in step (f). This conventional approach is computationally demanding due to the necessity of a large dataset in step (b). In this workflow, second-order force constants can be calculated in the conventional approach because the computational demand is much less than the calculation of third-order force constants in the conventional approach.
@@ III.1 LTC calculation
LTC values were computed by solving the Peierls-Boltzmann equation within the relaxation time approximation (RTA)[1, 2, 16] using the phono3py code.[9, 17] It is important to note that the method employed for the LTC calculations may not be appropriate for compounds with low or high thermal conductivity among the 103 compounds selected. To simplify the methodological investigation in this study, we considered only phonon-phonon scattering for determining the phonon relaxation times. Imaginary parts of phonon self-energies corresponding to the bubble diagram were calculated from supercell force constants, where a linear tetrahedron method was used to treat the delta function,[17] and their reciprocals were used as the relaxation times. Computational details on the calculation of the supercell force constants are provided in the subsequent sections. Phonon group velocities, mode heat capacities, frequencies, and eigenvectors were obtained from dynamical matrices constructed using second-order supercell force constants. Additionally, a non-analytical term correction [18, 19, 20] was applied to the dynamical matrices to account for long-range dipole-dipole interactions in the harmonic phonon calculation. For the wurtzite-type compounds, reciprocal spaces were sampled using a $19\times 19\times 10$ mesh, while a $19\times 19\times 19$ mesh was employed for the zincblende-type and rocksalt-type compounds.
@@ III.2 Supercell force constants calculation
The linear regression method was employed to calculate supercell force constants from the displacement-force dataset. In this method, forces acting on atoms, which were randomly displaced from their equilibrium positions in supercells, were computed either using the polynomial MLPs in our current approach or through first-principles calculations in the conventional approach. Tensor elements of supercell force constants were projected onto the subspace defined by symmetry projection operators of totally symmetric irreducible representations of the space group, index permutation, and translational invariance. In addition, detailed techniques were developed to enhance the efficiency of this computational process. This process is implemented in the symfc code.[21, 22] The crystallographic symmetries were determined using the spglib code.[23]
For the zincblende- and rocksalt-type compounds, we utilized supercells with $2\times 2\times 2$ and $4\times 4\times 4$ expansions of the conventional unit cells to calculate third-order and second-order supercell force constants, respectively. In the case of the wurtzite-type compounds, supercells with $3\times 3\times 2$ and $5\times 5\times 3$ expansions of the unit cells were employed for third-order and second-order supercell force constants calculations, respectively.
To compute the second-order supercell force constants, we employed the finite difference method as implemented in the phonopy code.[17, 15] We used the same displacement-force datasets as those in Ref. 15, where the forces in these datasets had been computed through first-principles calculations. The number of supercells in the datasets were six, two, and two for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively.
Third-order supercell force constants were calculated from the displacement-force datasets using the symfc code.[21] The numbers of symmetrically independent force constant elements were 7752, 1536, and 758 for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively. These values were determined based on the forces acting on atoms, which were inferred using the polynomial MLPs implemented in the pypolymlp code.[14]
@@ III.3 Polynomial MLPs
The polynomial MLPs were trained using the dataset composed of forces and displacements of atoms and energies in supercells. These energies and forces were computed through first-principles calculations. The performance of the polynomial MLPs in the LTC calculation via the third-order supercell force constants is discussed in Section IV.
For the 103 compounds, we trained the polynomial MLPs using the pypolymlp code.[14] In this training, Gaussian-type radial functions were employed, and the functional form $f_{n}(r)$ is given as
@@
where $r$ represents the distance from the center of each atom, and $r_{\text{c}}$ is the cutoff distance. $\beta_{n}$ and $r_{n}$ are the parameters, respectively. Radial functions with $r_{\text{c}}=8.0$ and $\beta_{n}=1.0$-2 and $r_{n}=(n-1)(r_{\text{c}}-1.0)/11$ for $n=1,\ldots,12$ were used. We considered polynomial invariants up to third order characterizing neighboring atomic density based on spherical harmonics with the maximum angular numbers of spherical harmonics $l_{\text{max}}^{(2)}=l_{\text{max}}^{(3)}=8$. The polynomial models were then constructed by the polynomial functions of the pair invariants and linear polynomial function of the polynomial invariants. We considered polynomial functions up to second order. The model coefficients were estimated from electronic total energies and forces by the linear ridge regression method.
@@ III.4 First-principles calculation
For the first-principles calculations, we employed the plane-wave basis projector augmented wave (PAW) method [24] within the framework of DFT as implemented in the VASP code.[25, 26, 27] The generalized gradient approximation (GGA) of Perdew, Burke, and Ernzerhof revised for solids (PBEsol) [28] was used as the exchange correlation potential. To ensure high numerical accuracy in computing atomic forces, the projection operators were applied in reciprocal spaces and additional support grids were employed for the evaluation of the augmentation charges. Static dielectric constants and Born effective charges were calculated with the conventional unit cells from density functional perturbation theory (DFPT) as implemented in the VASP code.[29, 30]
@@ III.5 Automation of dataset preparation
For the systematic calculations of energies and forces in supercells using first-principles calculations, we utilized the AiiDA environment [31, 32, 33] in conjunction with the AiiDA-VASP plugin.[34] The advantage of using the workflow automation system was not only the automation of submitting calculation jobs to high-performance computers, but also the automated data storing of the calculation results in a database systematically. The computed data, stored within the AiiDA database, could be conveniently accessed through the Python programming language. By writing a concise Python script, we were able to extract supercell energies, forces, and displacements from the AiiDA database on demand and convert this data into the structured format required for immediate use by the phono3py code.[9, 17] An example of the phono3py data format can be found in the phono3py github repository.
@@ III.6 Parameters for 103 binary compounds
33 compounds for the wurtzite- and zincblende-type and 37 compounds for the rocksalt-type were used to evaluate the LCT calculation approach proposed in this study, and their chemical compositions are listed in Tables 1 and 2. Crystal structures of the wurtzite and zincblende types are similar, though their stacking orders are different, much like the relationship between face-centered-cubic and hexagonal-close-packed structure types. Since it is of interest to explore their similarities and differences in calculations, as also studied in Ref. 9, the compounds with the same chemical compositions for the wurtzite and zincblende types were calculated. The tables also provide information on lattice parameters, the choices of PAW datasets from the VASP package, and electronic total energies of the elements that were subtracted from the total energies of the compounds used to train the polynomial MLPs.
Table 1: Lattice parameters, names of the VASP PAW-PBE datasets, electronic total energies of the atoms used in this study for 33 wurtzite- and zincblende-type compounds. w-$a$, w-$c$, and z-$a$ denote the lattice parameters $a$ and $c$ of the wurtzite-type compounds, and $a$ of the zincblende-type compounds, respectively.
@@ IV.2 Calculated LTCs
LTC values at 300 K predicted by the conventional approach, which directly uses the displacement-force dataset obtained through first-principles calculations to fit third-order supercell force constants, are depicted by the horizontal dotted lines. The third-order supercell force constants were computed by the linear regression method as implemented in the symfc code [21] from the first datasets with 100 supercells and 0.03 Å random directional displacements, which were those prepared for training the polynomial MLPs, as explained in Sec. IV.1. In addition, LTC values with 400 supercells for the zincblende- and rocksalt-type compounds and those with 400 and 2000 supercells for the wurtzite-type compounds were also computed. These values are depicted as horizontal lines in Figs. 2, 3, and 4. For most of the zincblende- and rocksalt-type compounds, LTC values derived from datasets with 100 supercells are found to be adequate when compared to those from 400 supercells. However, for the wurtzite-type compounds, even datasets with 400 supercells are insufficient.
The LTC values predicted for the wurtzite-type compounds using polynomial MLPs tend to align with those calculated directly from 2000 supercell datasets. This alignment emphasizes the utility and effectiveness of using polynomial MLPs in these cases.
Figure 2: Filled circles show LTCs ($\kappa$) of the 33 wurtzite-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The LTC values are the averages of the diagonal elements, i.e., $(2\kappa_{xx}+\kappa_{zz})/3$. The horizontal dotted, dashed-dotted, and dashed lines depict the LTC values calculated in the conventional approach from the datasets of 100, 400, and 2000 supercells without using the polynomial MLPs, respectively.
Figure 3: Filled circles show LTCs of the 33 zincblende-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The horizontal dotted and dashed-dotted lines depict the LTC values calculated in the conventional approach from the datasets of 100 and 400 supercells without using the polynomial MLPs, respectively.
Figure 4: Filled circles show LTCs of the 37 rocksalt-type compounds calculated at 300 K with respect to the number of supercells in the datasets used to train the polynomial MLPs. The horizontal dotted and dashed-dotted lines depict the LTC values calculated in the conventional approach from the datasets of 100 and 400 supercells without using the polynomial MLPs, respectively.